# Narrow line AGN selection in CEERS: spectroscopic selection, physical properties, X-ray and radio analysis

### v1

@@ 4.10 Radio stacking to discriminate the nature of X-ray weakness
To attempt to discriminate between the heavily obscured versus intrinsic X-ray weakness scenario to explain the observed X-ray weakness of these AGNs, we took into account the 1.4GHz radio image of the EGS field (Ivison et al., 2007). Different works showed that the physical processes and the magnetic fields at the origin of the hot corona, where the inverse Compton scattering produces the AGN X-ray emission, are also responsible for (at least part of) the synchrotron radiation at the origin of the AGN-related radio emission (Laor & Behar, 2008; Panessa et al., 2019). The connection between these two emissions is also at the basis of the so-called fundamental plane relations (Merloni et al., 2003) or of X-ray - radio luminosity relations reported in different works in the literature (D’Amato et al., 2022; Panessa et al., 2019; Fan & Bai, 2016).
@@
However, some works show that even the radio emission can be partially or completely absorbed by the BLR clouds if their uniform distribution is also dense enough to determine free-free absorption (Juodžbalis et al., 2024a, Mazzolari et al. in prep.). However, it is still not clear whether the X-ray and radio emissions originate and are distributed on the same scale, and, therefore, if the effect of obscuration from BLR clouds can suppress both or not (Paul et al., 2024). For example, in Laor & Behar (2008) the authors found that the smallest possible scale of radio emission () in radio-quiet (RQ) AGN, e.g., coming from a compact radio corona or jet base, should originate from a region times the extent of the X-ray emitting core. In this case, the radio emission would probably originate (or be being mainly distributed) outside of the BLR without being absorbed. Therefore in the following computations we assume that there is not a free-free absorption effect on the radio emission of the NLAGN and BLAGN investigated in this work, and we leave a more detailed discussion to Mazzolari et al. in prep.
@@
Consequently, if these AGN are normal AGN (in terms of bolometric to intrinsic X-ray luminosity), and heavily obscured (but without absorbing also the radio emission), we might expect not to detect their X-ray emission, but we would expect to detect their radio emission at the level returned by the intrinsic X-ray to radio luminosity relations derived in the literature. On the contrary, if the radio image is deep enough to detect the expected AGN radio emission, but none is detected, we would prefer a scenario in which both the radio and X-ray emissions are intrinsically lower with respect to what is observed in normal AGN at lower redshifts.
@@
None of the NLAGN or BLAGN are detected in the 1.4GHz image of the field, apart from two sources CEERS-2900 and CEERS-3129. Therefore, we perform a detailed stacking analysis at the position of the sources, as described in Sec. 2.5, considering the same three samples as for the X-ray stacking. The radio stacking analysis didn’t recover any 1.4GHz detection, but allowed us to derive upper limits on the median radio luminosities of these sources in the three samples. We compared these values with the expected AGN radio luminosities. Given the expected X-ray luminosities of the sources derived in Sec. 4.9, we can estimate the expected 1.4GHz luminosities using the luminosity relation derived for radio-quiet (RQ) AGN by D’Amato et al. (2022). The assumption that the sources are RQ is reasonable given that the fraction of radio-loud (RL) AGN has been found to be of the radio AGN population by different works, and it is not observed to vary with redshift (Liu et al., 2021; Williams & Röttgering, 2015).
@@
It is worth noting that the effectiveness of this test also depends on the fact that the AGN-related radio emission has to be dominant over the radio emission coming from star formation in the host galaxy. To check this, we considered the SFR derived for these AGNs from the SED fitting in Sec. 4.8, and we converted them into radio luminosities using the relations reported in Novak et al. (2017); Delvecchio et al. (2021). For all the three AGN samples the median 1.4GHz luminosities due to star formation are 1.5 dex lower than the AGN radio luminosity derived from the luminosity relation of D’Amato et al. (2022).
@@
The comparison between the observed and the expected 1.4GHz luminosities of the three AGN samples is presented in Fig. 16, together with the upper limits of the single sources in the three samples. Unfortunately, the AEGIS20 radio image is too shallow to give conclusive constraints on the nature of the X-ray weakness of these samples of AGN, being the upper limits of the stack still compatible with the expected radio luminosities. We also explore the possibility that the bulk of the sources are instead RL, and, therefore, we computed the expected radio luminosities of the three stacks considering the luminosity relation of Fan & Bai (2016), derived for RL AGN. In this case, the upper limits on the 1.4GHz luminosities are different orders of magnitudes lower than expected, excluding the RL scenario, which was, however, already considered improbable implying all (or most of) the sources to be RL.
@@
In Fig. 16, we also report the position of the two radio-detected sources, i.e., CEERS-2900 and CEERS-3129, at and , respectively. The first has a 1.4GHz radio flux of 58mJy, which corresponds to an observed radio luminosity of erg/s. If we assume that CEERS-2900 is a RQ AGN, the expected radio luminosity given its intrinsic 2-10keV X-ray luminosity reported in Buchner et al. (2015) (where the source is classified as a CTK AGN) is only erg/s. This strongly points toward the RL nature of this source. Indeed, considering the RL luminosity relation reported in Fan & Bai (2016), the observed and expected radio luminosities correspond almost perfectly.
@@
The value of the expected , computed starting from the expected luminosity, is almost consistent with the hypothesis of CEERS-3129 being a RQ AGN, since the observed and expected radio luminosities are quite consistent (close to the scatter of the relation). This is even more true if we recall that the expected , derived from the dust-corrected computed using the broad H emission, is probably slightly underestimated because the obscuration of the broad H component is probably larger than the one derived from the narrow-line Balmer decrement (see Sec. 4.9).
The fact that the expected and observed radio luminosities of CEERS-3129 almost correspond supports the extreme CTK scenario to explain the X-ray weakness of this source. However, this scenario is not free from problematics. For example, it requires the CTK envelope around the central SMBH absorbing the X-ray emission to be almost completely dust-free to allow the broad H emission to escape. A natural solution to this problem is that the BLR clouds themselves, with large column densities distributed with approximately a covering factor, absorb the X-ray photons, as discussed and analyzed in Maiolino et al. (2024a).
@@ 5 Conclusions
In this work, we analyzed the public JWST/NIRSpec medium resolution (MR) spectra of the CEERS survey (PID.1340) to identify NLAGN using multiple emission-line diagnostic diagrams. In particular, we identify a secured spectroscopic redshift of 217/313 targets observed with MR setup. We performed spectral line fitting using QubeSpec, and fitting separately different complexes of emission lines, ranging from rest-UV to rest-optical. We considered 10 different emission line diagnostic diagrams to perform the NLAGN selection, using both demarcation lines taken from the literature and also exploiting the distribution of SFG and AGN photoionization models taken from Gutkin et al. (2016) and Feltre et al. (2016). This process led to the identification of 52 NLAGN, whose IDs and main physical parameters are reported in Table 3 to allow the scientific community to perform further studies. The main results are the following:
@@
Traditional AGN diagnostic diagrams, in particular the BPT (R3N2), are less effective in selecting AGN at , because high- galaxies and AGN generally overlap in these diagnostics due to the different physical conditions of high- sources. For the high- regime, the most effective AGN diagnostic diagrams are those based on the [O iii] 4363 line or on the high-ionization emission lines from Mazzolari et al. (2024).
@@
We select 10 NLAGN at (3 of them marked as candidates NLAGN), whose spectra are reported in Appendix D.
@@
The NLAGN selection includes only one out of the eight BLAGN at selected in (Harikane et al., 2023), i.e. CEERS-1244, four out of five of the X-ray sources with a CEERS MR spectrum and classified as AGN by the X-ray spectral analysis in Buchner et al. (2015), and both the radio detected sources reported in the radio catalog of Ivison et al. (2007).
@@
The fraction of NLAGN selected among the CEERS spectra analyzed in this work is , a fraction that is almost constant across different redshifts and agrees well with what was found by Scholtz et al. (2023a) for the JADES survey.
@@
We analyzed the distribution of [O iii] 5007 NL FWHM of the AGN and non-AGN samples without finding any significantly different distribution for the two populations, suggesting that the NLAGN identified in this work do not significantly impact the ISM of their host galaxies. This probably happens because, thanks to the sensitivity of JWST, we are sampling a different population of NLAGN without outflows compared to previous NLAGN samples. Furthermore, in many cases, and in particular for low-mass high-redshift sources, the FWHM are very close to the resolution limit of JWST.
@@
We also investigated the average obscuration level of the sources by computing the values for those sources having both the H and H lines available in their spectra. We did not find a significant evolution with a redshift of the value of . We also note that the AGN population is characterized, in the lower redshift bin , by a mag larger obscuration value, while in the other redshift bins the median value of obscuration of NLAGN and SFG is almost the same.
@@
We computed the bolometric luminosities of the sources using the dust-corrected H fluxes and the scaling relation reported in Netzer (2009). The range of spanned by our NLAGN sample at is much lower than the AGN bolometric luminosities reported in any pre-JWST samples, but dex larger than bolometric luminosities of the sample of JWST detected NLAGN reported in Scholtz et al. (2023a) selected in the JADES survey (where the observing times are significantly longer). We tested the robustness of the derived bolometric luminosities by comparing our baseline results with those derived using the calibrations reported in Lamastra et al. (2009) for the [O iii] 5007 line, finding a good agreement between the two.
@@
We performed a SED-fitting analysis using CIGALE to derive the main physical properties of the host galaxies of the sample NLAGN and non-AGN. We found that, on average, sources at deviate from the main sequence of Popesso et al. (2023), moving toward higher values of SFR than what is predicted by the MS. The range of stellar masses extends down to . We didn’t find any significant difference in the SFR or distribution of the NLAGN compared to those of the parent sample, suggesting that any potential AGN positive or negative effect on the host-galaxy SFR is not significantly taking place. This is further supported by the lack of significant outflow detections in their spectra.
@@
We investigate the X-ray properties of the selected NLAGN and of the BLAGN selected in Harikane et al. (2023). We found that none of the BLAGN is detected in the X-ray image of the field and that only 4 of the 52 NLAGN are X-ray detected. We performed an X-ray stacking analysis on the AGN that are not X-ray detected, finding an X-ray deficit of dex for the high- BLAGN and for the sample of NLAGN at , and dex for the NLAGN at between their expected X-ray luminosities (given their bolometric luminosity) and what comes from the X-ray stacking. We tried to investigate the origin of this X-ray weakness by studying the radio emission of these sources, which can help to disentangle the heavily obscured AGN scenario from the intrinsic X-ray weak one. Unfortunately, the radio observation covering the region of the CEERS survey is not deep enough to provide a conclusive answer. However, this shows the key role that present and future deep radio continuum observations can have in understanding the problem of the X-ray weakness of high- AGN.
@@
Interestingly, one of the low-redshift BLAGN, which has also been selected by the NLAGN selection, CEERS-3129, is undetected in the deep X-ray image (showing an X-ray deficit of dex), but it is detected in the radio image of the field, with a radio luminosity compatible with the RQ AGN nature. The radio detection of CEERS-3129 supports the CTK AGN scenario at the origin of the observed X-ray weakness.
@@
JWST is allowing us to reveal the population of heavily obscured AGN up to very high redshift () and, at the same time, to investigate the NLAGN population at low redshift with a spectroscopic sensitivity not explored by pre-JWST surveys. This work highlights the necessity of defining new NLAGN diagnostic diagrams and, in general, different and complementary AGN selection techniques suited for the identification at high redshift AGN. Future works will have to explore in more detail the actual number density and the demography of these objects in the early Universe, carefully accounting for the selection function of each spectroscopic campaign. We also stress that a multiwavelength approach, such as the one outlined in this work, is necessary to explore and potentially provide answers to the peculiar features characterizing high redshift AGN.

### v3

@@ 4.10 Radio stacking to discriminate the nature of X-ray weakness
To discriminate between the heavily obscured versus intrinsic X-ray weakness scenario and explain the observed X-ray weakness of these AGNs, we took into account the 1.4GHz radio image of the EGS field (Ivison07). Different works have shown that the physical processes and the magnetic fields at the origin of the hot corona, where the inverse Compton scattering produces the AGN X-ray emission, are also responsible for (at least part of) the synchrotron radiation at the origin of the AGN-related radio emission (Laor08; panessa19). The connection between these two emissions is also at the basis of the so-called fundamental plane relations (merloni03) or of X-ray - radio luminosity relations reported in different works in the literature (damato22; panessa19; fan16).
@@
However, some works show that even the radio emission can be partially or completely absorbed by the BLR clouds if their uniform distribution is also dense enough to determine free-free absorption (Juodzbalis24_rosetta, Mazzolari et al. in prep.). However, it is still not clear whether the X-ray and radio emissions originate and are distributed on the same scale, and, therefore, if the effect of obscuration from BLR clouds can suppress both or not (Paul24). For example, in Laor08 the authors found that the smallest possible scale of radio emission () in radio-quiet (RQ) AGN, e.g., coming from a compact radio corona or jet base, should originate from a region times the extent of the X-ray emitting core. In this case, the radio emission would probably originate (or be being mainly distributed) outside of the BLR without being absorbed. Therefore, in the following, we assume that there is not a free-free absorption effect on the radio emission of the NLAGN and BLAGN investigated in this work, and we leave a more detailed discussion to Mazzolari et al. in prep.
@@
Consequently, if these AGNs are normal AGNs (in terms of bolometric to intrinsic X-ray luminosity), and heavily obscured (but without absorbing also the radio emission), we might expect not to detect their X-ray emission, but we would expect to detect their radio emission at the level returned by the intrinsic X-ray to radio luminosity relations derived in the literature. On the contrary, if the radio image is deep enough to detect the expected AGN radio emission, but none is detected, we would prefer a scenario in which both the radio and X-ray emissions are intrinsically lower with respect to what is observed in normal AGN at lower redshifts.
@@
None of the NLAGNs or BLAGNs are detected in the 1.4GHz image of the field, apart from two sources CEERS-2900 and CEERS-3129. Therefore, we performed a detailed stacking analysis at the position of the sources, as is described in Sec. 2.5, considering the same three samples as for the X-ray stacking. The radio stacking analysis did not recover any 1.4GHz detection, but allowed us to derive upper limits on the median radio luminosities of the sources in the three samples. We compared these values with the expected AGN radio luminosities. Given the expected (and intrinsic) X-ray luminosities of the sources derived in Sec. 4.9, we can estimate the expected 1.4GHz luminosities using the luminosity relation derived for radiatively efficient radio-quiet (RQ) AGN by damato22. The assumption that the sources are RQ is reasonable given that the fraction of radio-loud (RL) AGN has been found to be of the radio AGN population by different works, and it is not observed to vary with redshift (liu21; williams+15).
@@
It is worth noting that the effectiveness of this test also depends on the fact that the AGN-related radio emission has to be dominant over the radio emission coming from star formation in the host galaxy. To check this, we considered the SFRs derived for these AGNs from the SED fitting in Sec. 4.8, and we converted them into radio luminosities using the relations reported in novak17; delvecchio21. For all the three AGN samples the median 1.4GHz luminosities due to star formation are 1.5 dex lower than the AGN radio luminosity derived from the luminosity relation of damato22.
@@
The comparison between the observed and the expected 1.4GHz luminosities of the three AGN samples is presented in Fig. 16, together with the upper limits of the single sources in the three samples. Unfortunately, the AEGIS20 radio image is too shallow to give conclusive constraints on the nature of the X-ray weakness of these AGN, being the upper limits of the stack compatible with the expected radio luminosities. We also explore the possibility that the bulk of the sources are instead RL, and, therefore, we computed the expected radio luminosities of the three stacks considering the luminosity relation of fan16, derived for radiatively inefficient RL AGN. In this case, the observed upper limits on the 1.4GHz luminosities are different orders of magnitudes lower than expected. This excludes that the bulk of our sources is dominated by RL AGN.
@@
In Fig. 16, we also report the position of the two radio-detected sources, CEERS-2900 and CEERS-3129, at and , respectively. The first has a 1.4 GHz radio flux of 58mJy, which corresponds to an observed radio luminosity of erg/s. If we assume that CEERS-2900 is a RQ AGN, the expected radio luminosity given its intrinsic 2-10keV X-ray luminosity reported in buchner15 (where the source is classified as a CTK AGN) is only erg/s. This strongly points toward the RL nature of this source. Indeed, considering the RL luminosity relation reported in fan16, the observed and expected radio luminosities correspond almost perfectly.
@@
The value of the expected , computed starting from the expected luminosity, is almost consistent with the hypothesis of CEERS-3129 being a RQ AGN, since the observed and expected radio luminosities are consistent within 2. This is even truer if we recall that the expected , derived from computed using the dust-corrected broad H emission, is probably underestimated because the obscuration of the broad H component is probably larger than the one derived from the narrow-line Balmer decrement (see Sec. 4.9).
The fact that the expected and observed radio luminosities of CEERS-3129 almost correspond supports the extreme CTK scenario to explain the observed X-ray weakness of this source. However, this scenario is not free from problematics. For example, it requires the CTK envelope around the central SMBH absorbing the X-ray emission to be almost completely dust-free to allow the broad H emission to escape. A natural solution to this problem is that the BLR clouds themselves, with large column densities distributed with approximately a covering factor, absorb the X-ray photons, as discussed and analyzed in Maiolino24_X.
@@ 5 Conclusions
In this work, we have analyzed the public JWST/NIRSpec MR spectra of the CEERS survey (PID.1340) to identify NLAGNs using multiple emission-line diagnostic diagrams. First, we attributed spectroscopic redshift to 217 out of the 313 targets observed with the JWST/NIRSpec MR setup. We performed spectral line fitting using QubeSpec, and fitting separately different complexes of emission lines, ranging from rest-UV to rest-optical. We considered ten different emission line diagnostic diagrams to perform the NLAGN selection, using both demarcation lines taken from the literature and also exploiting the distribution of SFG and AGN photoionization models taken from Gutkin16 and Feltre16. This process led to the identification of 52 NLAGNs, whose IDs and main physical parameters are reported in Table 8 to allow the scientific community to perform further studies. The main results are the following:
@@
Traditional AGN diagnostic diagrams, in particular the BPT (R3N2), are less effective in selecting AGNs at , because high- galaxies and AGNs generally overlap in these diagnostics due to the different physical conditions of high- sources. For the high- regime, the most effective AGN diagnostic diagrams are the ones based on the [O iii] 4363 line (Mazzolari24b) or on the high-ionization emission lines.
@@
We selected ten NLAGNs at (four of them marked only as candidate NLAGNs), whose spectra are reported in Appendix D.
@@
The NLAGN selection includes only one out of the eight BLAGNs at selected in (Harikane23) – CEERS-1244 – four out of five of the X-ray sources with a CEERS MR spectrum and classified as AGNs by the X-ray spectral analysis in buchner15, and both the radio detected sources reported in the radio catalog of Ivison07.
@@
NLAGN represent of the sources analyzed in this work. This fraction is almost constant across different redshifts and agrees well with what was found by Scholtz23b for the NLAGN population selected among the spectra of the JADES survey.
@@
We analyzed the distributions of [O iii] 5007 narrow-line FWHMs of the AGN and non-AGN samples without finding significant differences in the distributions of the two populations. This suggests that the NLAGNs identified in this work are not significantly impacting the ISM of their host galaxies. This is probably due to the fact that, thanks to the sensitivity of JWST, we are selecting a fainter NLAGN population that is characterized by having a weaker impact on the host galaxy and lacking powerful outflows (contrary to what has been observed for other NLAGN samples at lower redshifts). Furthermore, in many cases, and in particular for low-mass high-redshift sources, the FWHMs are very close to the resolution limit of JWST.
@@
We also investigated the average obscuration level of the sources by computing the values from the Balmer decrements. We did not find a significant evolution with redshift in the value of . We also note that the AGN population is characterized, in the lower-redshift bin , by a mag larger obscuration, while in the other redshift bins the median value of obscuration of NLAGNs and non-AGN sources is almost the same.
@@
We computed the bolometric luminosities of the sources using the dust-corrected H fluxes and the scaling relation reported in Netzer09. The range of spanned by our NLAGN sample at is much lower than the AGN bolometric luminosities reported in any pre-JWST samples, but dex larger than bolometric luminosities of the JWST selected NLAGN reported in Scholtz23b selected in the JADES survey (where the exposure times are significantly longer). We tested the robustness of the derived bolometric luminosities by comparing our baseline results with the ones derived using the calibrations reported in Lamastra09 for the [O iii] 5007 line, finding a good agreement between the two.
@@
We performed a SED-fitting analysis using CIGALE to derive the main physical properties of the host galaxies for the sample of NLAGNs and non-AGNs. We found that, on average, sources at deviate from the main sequence of Popesso23, moving toward higher SFR values than what is predicted by the MS. The range of stellar masses extends down to . We did not find any significant difference in the SFR or distribution of the NLAGN compared to the ones of the parent sample, suggesting that any potential positive or negative AGN feedback on the host-galaxy SFR is not significantly taking place. This is in agreement with the lack of significant outflow detections in their spectra.
@@
We investigated the X-ray properties of the selected NLAGNs and of the BLAGNs selected in Harikane23. We found that none of the BLAGNs is detected in the X-ray image of the field and that only 4 of the 52 NLAGNs are X-ray detected. We performed an X-ray stacking analysis on the AGNs that are not X-ray-detected, finding a significant X-ray deficit between their expected X-ray luminosities (given their bolometric luminosity) and what comes from the X-ray stacking. This observed X-ray weakness is dex for the high- BLAGNs and for the sample of NLAGN at , and dex for the NLAGN at . We tried to investigate the origin of the X-ray weakness by studying the radio emission of these sources, which can help to disentangle the heavily obscured AGN scenario from the intrinsic X-ray weak one. Unfortunately, the radio observation covering the region of the CEERS survey is not deep enough to provide a conclusive answer. However, this shows the key role that present and future deep radio continuum observations can have in helping us understand the problem of the X-ray weakness of high- AGNs.
@@
Interestingly, one of the low-redshift BLAGNs, CEERS-3129, is undetected in the deep X-ray image (showing an X-ray deficit of dex), but it is detected in the radio image of the field, with a radio luminosity compatible with a RQ AGN nature. The radio detection of CEERS-3129 supports the CTK AGN scenario at the origin of the observed X-ray weakness.
@@
JWST is allowing us to reveal the population of heavily obscured AGNs up to very high redshifts () and, at the same time, to investigate the NLAGN population at low redshifts with a spectroscopic sensitivity never explored by pre-JWST surveys. This work highlights the necessity of defining new NLAGN diagnostic diagrams and, in general, different and complementary AGN selection techniques suited for the identification of high-redshift AGNs. Future works will have to explore in more detail the actual number density and the demography of these objects in the early Universe, carefully accounting for the selection function of each spectroscopic campaign. We also stress that a multiwavelength approach, such as the one outlined in this work, is necessary to explore and potentially provide answers to the peculiar features characterizing high-redshift AGNs. In this sense, the future Square Kilometer Array Observatory (SKAO) and the Advanced X-Ray Imaging Satellite (AXIS) will be crucial to reveal the actual radio and X-ray properties of these objects.